OpenStudy (anonymous):

What is the amplitude and period of :H(x)=-3sin(x- pi/4) +2, with respect to the parent function F(x)=sinx?

hartnn (hartnn):

hint : the amplitude of Y(x) = Asin (x+theta)+B is |A|+B and its phase shift is theta.

hartnn (hartnn):

and period remains unchanged with any phase shift and is independent of amplitude. so period of H(x)=-3sin(x- pi/4) +2 is same as that of sin x, which is ..... ?

OpenStudy (anonymous):

-3?

hartnn (hartnn):

period ? no period of sin x is 2\(\pi\).

OpenStudy (anonymous):

so the period for the equation is 2pi?

hartnn (hartnn):

yes. and u got the amplitude ?

OpenStudy (anonymous):

no, How do I find that?

hartnn (hartnn):

Y(x) = Asin (x+theta)+B is |A|+B for your question A=-3B=2

hartnn (hartnn):

A=-3, B=2

OpenStudy (anonymous):

so its -1?

hartnn (hartnn):

nopes, |-3| = 3

OpenStudy (anonymous):

so 4

OpenStudy (anonymous):

and the phase shift is pi/4?

hartnn (hartnn):

3+2 is 4 ??

OpenStudy (anonymous):

i mean 5 my bad

hartnn (hartnn):

phase shift is -pi/4

OpenStudy (anonymous):

oh ok thanks!

hartnn (hartnn):

welcome ^_^